OpenStudy (anonymous):

Graph using intercepts -2x+3y=6

OpenStudy (anonymous):

to find the \(y\) intercept, put \(x=0\) and solve for \(y\) you get \[3y=6\iff y=2\] so the \(y\) intercept is \((0,2)\) for the \(x\) intercept, put \(y=0\) and solve \[-2x=6\]

OpenStudy (amistre64):

divide by 6, then divide off the coeffs: the under parts define the intercepts .. its called the intercept form

OpenStudy (amistre64):

\[-2x+3y=6\] \[\frac{-2x}{6}+\frac{3y}{6}=1\] \[\frac{x}{-6/2}+\frac{y}{6/3}=1\]